OpenStudy (anonymous):

solve the linear system by graphing x+3y=2 and 3x+9y=6

OpenStudy (anonymous):

Not by graphing oops

OpenStudy (anonymous):

Answer is 2-3k, k

OpenStudy (anonymous):

Each one is the equation of a straight line. x+3y=2 3x-9y=6 multiply first by 3 and add 3x+9y=6 3x–9y=6 6x = 12 x = 2 2+3y=2 y = 0 so they are two lines intersecting at x=2 and y=0

OpenStudy (anonymous):

so, 2-3k = 2 and k = 0

OpenStudy (anonymous):

They are both addition equations

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

can u please vote http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

There is no negative 9

OpenStudy (anonymous):

yeah, thats incorrect x+3y = 2 3x + 9y = 6 so x+3y = 2 -x - 3y = -6 0 = -6 no solutions

OpenStudy (anonymous):

The answer is 2-3k,k

OpenStudy (anonymous):

Where k is a set of real numbers

OpenStudy (anonymous):

okay, then x = 2 - 3y 3(2-3y) +9y = 6 6 - 9y + 9y = 6 6 = 6 All real numbers. If you graph them they are the same line

OpenStudy (anonymous):

Thanks!